OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 please help :(

jimthompson5910 (jim_thompson5910):

Solve for y and get into y = mx + b form \[\Large -11x + 9y = -12 \] \[\Large 9y = -12 + 11x \] \[\Large 9y = 11x-12 \] \[\Large y = \frac{11x-12}{9} \] \[\Large y = \frac{11x}{9}-\frac{12}{9} \] \[\Large y = \frac{11}{9}x-\frac{4}{3} \] The equation is now in y = mx+b form where m = 11/9 b = -4/3

jimthompson5910 (jim_thompson5910):

m is the slope, so the slope is 11/9 b is the y-intercept, so the y-intercept is -4/3, which is the point (0, -4/3)